The sum of two integers is 25. The difference of their squares is 275. Find the two integers.
Answer by venugopalramana(3286)

let the 2 integers be x and y
we are given ....x+y=25......(1)... and x^2-y^2=275.
we know (x+y)(x-y)=x^2-y^2
so ...25(x-y)=275
x-y=275/25=11......(2)
(1)+(2)....x+y+x-y=25+11=36
2x=36
x=18..so ...18+y=25..or y=25-18=7